Hello,
I have many problems graphing equations such as: y=3X=8 How do you know where to graph it(put the dots and lines)? (without using a calculator) and solving systems if of equations by graphing
ex)
3x+y + -2
6x + 2y = 10

Answer by Alan3354(69443)

Hello,
I have many problems graphing equations such as: y=3X=8 How do you know where to graph it(put the dots and lines)? (without using a calculator) and solving systems if of equations by graphing
ex)
3x+y + -2
6x + 2y = 10
--------------------
These are 2 linear equations. Linear means their graphs will be a straight line. To graph a straight line, you need 2 points, then draw the line thru the 2 points.
When there are 2 lines, you can see the "answer" where the lines cross, which is where x and y are the same for both equations.
---------
To graph 3x+y = 2 (I'll assume the second + was supposed to be an equal sign), find 2 points. One way to do that is to set x = 0 and solve for y, then set y = 0 and solve for x:
3x+y = -2
@ x=0: 3*0 + y = -2
y = -2, so the point is (0,-2)
@ y=0: 3x = -2
x = -2/3, so the point is (-2/3,0)
Draw a straight line thru the 2 points.
-------------
For the 2nd: 6x + 2y = 10
3x+y = 5
@ x=0: y = 5 --> (0,5)
@ y=0: 3x = 5
x = 5/3 --> (5/3,0)
Draw a straight line thru those 2 points.
The intersection of the 2 lines (if they intersect) is the "solution" to the pair of eqns, or the point where x and y fit both equations.